# LongViTU: Instruction Tuning for Long-Form Video Understanding

### v1

@@ 3.2 Quantitative Analysis on LongViTU
Frame-Based mPLUG-OWL 42.4 33.5 37.6 43.6 35.4 37.8
Video-LLaVA 45.9 37.8 46.3 49.1 38.1 42.7
@@
Frame-Based mPLUG-OWL 42.4 45.8 47.7 47.7 49.4 47.6
Video-LLaVA 45.9 45.6 50.5 48.8 53.2 49.4
@@
Frame-Based mPLUG-OWL 42.4 42.1 54.6 54.3 51.5 50.3
Video-LLaVA 45.9 41.6 56.8 55.3 54.6 51.7
@@
LLaVA-OneVision 60.1 58.2 69.1 53.3 46.7 - - - -
LongVU 67.6 56.2 66.1 54.7 47.9 35.7 48.3 51.1 42.8
LongVU∗∗ 69.1 56.3 65.8 54.7 48.4 36.5 48.9 51.4 44.2
@@
Spatial vs. Temporal Bias.  The weaker performance in Spatiotemporal Understanding reveals the unresolved challenges in long-form video comprehension. Tasks requiring detailed spatial analysis are particularly demanding. In contrast, Episodic Reasoning and Commonsense Inference leverage the logical sequence of events, resulting in improved performance through text-based methods.
Limitations of Commercial Models.  The commercial model Gemini-1.5-Pro demonstrates unexpectedly poor performance, marginally surpassing the zero-shot capabilities of open-source models while falling behind the fine-tuned open-source model LongVU. Due to resource constraints, full-length videos with specific timestamps embedded in the prompts were used instead of segmenting videos for each QA pair. Consequently, Gemini-1.5-Pro was required to process videos ranging from tens of minutes to over two hours for many questions, which significantly degraded its performance. These findings highlight the substantial challenges inherent in long-form video understanding.
3.3 Quantitative Analysis on Benchmarks
This section evaluates performance on benchmarks covering both In-Distribution (ID) and Out-of-Distribution (OOD) scenarios, leveraging datasets EgoSchema [36], VideoMME [16], WorldQA [67], and OpenEQA [35], as detailed in subsection 3.2. Key findings are outlined below.
Limitations of Sampling Models.  The LLaMA-VID [30] model exhibited a substantial performance drop during zero-shot evaluation in stage 3, which focuses on fine-tuning for long videos, compared to its performance in stage 2, which involves pre-training on images and short videos. This disparity underscores critical deficiencies in the stage 3 fine-tuning strategy. We revised the fine-tuning approach by applying all adjustments used in stage 2 as the baseline. The results demonstrated significant improvements across all benchmarks, with a maximum gain of 11.8% on EgoSchema.
Better on Longer Videos.  Fine-tuning with LongViTU resulted in decreased performance on shorter videos, as observed in the VideoMME Short subset and OpenEQA results. The average video duration of OpenEQA (49 seconds) is even shorter than that of the VideoMME Short subset (83 seconds). In contrast, performance on the Medium (563 seconds) and Long (2386 seconds) subsets improved. Specifically, LongViTU fine-tuning led to a 1% improvement on the VideoMME Long subset, while performance on the Short subset decreased and the Medium subset remained stable. This highlights the effectiveness of LongViTU in enhancing comprehension for longer videos, albeit with some trade-offs in performance on shorter ones.
Long-form Video Understanding Challenge.  The VideoLLM Online [7] model struggled significantly with the Long subset of VideoMME, producing incoherent responses and failing to generate measurable evaluations, underscoring the challenges of understanding long-form videos. The LongViTU dataset serves as both a vital pre-training resource and a benchmark for advancing long-form video understanding. This is demonstrated by the improved performance of models like LongVU on various additional benchmarks after supervised fine-tuning (SFT) on LongViTU, emphasizing the importance of optimizing models with large-scale high-quality video question-answering data.
3.4 Qualitative Evaluation
We present visualizations of various question-answering types in Figure 7 to facilitate a more thorough qualitative analysis. Due to space constraints, these figures and the ablation study have been deferred to Appendix B.
Spatial Details.  In Figure 7(a), the dense arrangement of multiple foreground objects in the scene led to incorrect zero-shot predictions by both Video-LLaVA and LongVU. Fine-tuning with LongViTU enabled the model to focus on finer spatial features, resulting in correct answers.
Key Moments.  As illustrated in Figure 7(b), the fine-tuned models accurately identified the key moment, a brief appearance of "a plant on the windowsill", and provided a precise and concise response, achieving a perfect GPT-4 score of 100. In contrast, the Video-LLaVA zero-shot model failed to capture this brief key moment.
Temporal Localization.  In Figure 7(c), both Video-LLaVA and LongVU correctly recognized "two" plug-in sockets in the kitchen at the end of a lengthy video and provided accurate responses after fine-tuning. However, both models failed in the zero-shot setting. This highlights the difficulty of extracting temporal details from extended video sequences and emphasizes the contribution of LongViTU data in improving generalization for long-form temporal localization.
3.5 Future Research
LongViTU introduces a novel direction in streaming video processing and video memory storage mechanisms for end-to-end models. Future research will build on these contributions, aiming to integrate streaming video understanding with question-answering, thereby paving the way for a more holistic approach to video understanding.
4 Related Work
Large Language Models.  Large language models (LLMs), such as InstructGPT [38], GPT-4 [1], LLaMA [48], and LLaMA-2 [49], have demonstrated notable capabilities in text processing and have been applied to the generation of large-scale datasets. These models convert different modalities into structured textual descriptions, which can then be used to prompt GPT-4 to produce multimodal data.
Instruction Tuning.  LLaVA [33] was among the first to leverage foundational vision models to generate image captions and detect bounding boxes, which were subsequently processed by ChatGPT or GPT-4 to generate image-based QA pairs. Building on this, methods like Bongard-OpenWorld [53], Video-LLaVA [31], and VideoChat [28] extended these principles to a more complex pipeline, shifting the focus from individual image QA to video-based QA. By sampling multiple frames from videos and applying LLaVA’s procedure, these approaches generate video QA datasets using structured frame descriptions, object categories, and attributes. However, the repetitive nature of frame-based descriptions, coupled with the input length limitations of LLMs, constrains the number of frames, thereby reducing the comprehensiveness of datasets.
Long-Context Language Models.  Even the most advanced long-context LLMs [46], such as GPT-4-turbo, ChatGLM [18], Baichuan2 [57], and InternLM2 [6], capable of handling input sequences beyond 128k tokens, suffer significant performance degradation when dealing with long and intricate texts. They struggle to manage the redundancy and disorganization in detailed video descriptions, limiting their effectiveness in generating video QA. Unlike static images, videos inherently demand an understanding of temporal dynamics, making event correlation crucial for video comprehension. The current frame-based approach fails to address this temporal aspect adequately, often resulting in shallow QA generation confined to individual frame analysis.
Long-Form Video Understanding.  Instruction tuning in the LLaVA paradigm [33] has demonstrated significant potential for multimodal tasks like captioning [8] and visual question-answering [4, 3, 47, 27, 12, 58, 2, 20, 39, 52]. Despite their effectiveness for images and short videos, these approaches encounter challenges with long-form videos [45, 31, 34, 63, 50, 65]. The primary limitation arises from the large number of visual tokens produced by encoders, which can range from 576 to 2880 tokens per image in models like LLaVA-NeXT [32]. As frame counts increase, these tokens rapidly surpass the context window limits of LLMs. To address this, recent methods use resamplers to reduce the visual tokens before passing them to LLMs [27, 30, 5, 10]. However, such reductions often degrade visual representations, resulting in diminished performance. Alternative approaches, including advanced pruning and feature merging, offer promising solutions [9, 43, 23, 68]. Notably, recent models such as VideoLLM Online [7], LongVILA [56], LLaVA-OV [26], LongVA [65], and LongVU [44] signify substantial advancements in long-form video understanding.
@@ 5 Conclusion
We present LongViTU, a large-scale dataset designed for long-form video understanding with integrated video memory and explicit timestamp annotations. Its hierarchical tree structure efficiently organizes video content, addressing critical challenges in generating QA pairs for extended videos. The self-revision process further enhances QA quality by mitigating hallucinations, redundancy, and irrelevant details. Fine-tuning on LongViTU demonstrates significant performance gains on both LongViTU and additional benchmarks, confirming its effectiveness and generalizability. Future work will explore advanced memory storage mechanisms that leverage explicit timestamps for advancing long-form video understanding and streaming question-answering.
@@
Appendix A Limitation Statement
This paper focuses on the challenges of long-form video understanding and memory-based streaming video question-answering. Our automated pipeline demonstrates adaptability across diverse scenarios, with self-revision mechanisms that enhance data quality. However, some generated data may lack meaningful context, potentially impacting model performance. To mitigate this disadvantage, we propose refining the pipeline and incorporating manual inspection to better align the benchmark with real-world human annotations, and our work remains ongoing.
@@ Appendix B Additional Results
Due to space limitations, we defer more details about the ablation study, human study, and qualitative evaluation to this section.
@@ B.1 Ablation Study
Varying Durations.  Figure Figure 4 illustrates LongVU’s performance on LongViTU with varying QA durations. As the QA duration (in seconds) increases, the performance gradually declines, consistent with the trends observed in VideoMME across all baseline models in subsection 3.2.
Figure 4: Duration Ablation.  After fine-tuned on LongViTU, LongVU’s performance exhibits a mitigated downward trend on longer subsets [300, 600) compared to shorter subsets.
Training Epochs.  Figure Figure 5 illustrates LongVU’s performance on Egoschema improves after SFT as the training epochs derived from LongViTU increase. The performance gains diminish as the training epochs approach the full scale of LongViTU, eventually plateauing at its optimal performance.
@@
B.2 Human Study
We randomly selected 100 QA pairs from both LongViTU and VideoMME for a human study conducted by three evaluators, as shown in Figure 6.
Self-Evaluation.  Each QA pair was assessed for its alignment with the video content. Good pairs were concise, accurate, and spatiotemporally relevant with video content. Fair pairs matched the video but lacked spatiotemporal relevance. Poor pairs included irrelevant or inaccurate information.
LongViTU vs. VideoMME.  QA pairs were assessed for coherence, and alignment with video content, comparing their similarity to human-crafted pairs. The results were categorized as: VideoMME significantly better, LongViTU significantly better, or Tie for comparable quality.
Figure 6: Human Preference.  In self-evaluation, approximately 91% of QA pairs were marked as correct. In comparison, around 64% of QA pairs were comparable in quality to VideoMME.
B.3 Qualitative Evaluation
A detailed qualitative analysis is provided in subsection 3.4 and illustrated in Figure 7.
(a) Spatial Details.  The dense arrangement of multiple foreground objects in the scene led to incorrect zero-shot predictions by both Video-LLaVA and LongVU. Fine-tuning with LongViTU enabled the model to focus on finer spatial features, resulting in correct answers.
(b) Key Moments.  The fine-tuned models accurately identified the key moment, a brief appearance of "a plant on the windowsill", and provided a precise and concise response, achieving a perfect GPT-4 score of 100. In contrast, the Video-LLaVA zero-shot model failed to capture this brief key moment.
(c) Temporal Localization.  Both Video-LLaVA and LongVU correctly recognized "two" plug-in sockets in the kitchen at the end of a lengthy video and provided accurate responses after fine-tuning. However, both models failed in the zero-shot setting. This highlights the difficulty of extracting temporal details from extended video sequences and emphasizes the contribution of LongViTU data in improving generalization for long-form temporal localization.
@@ C.2 Long-Form QA Generation
QA Generation.  We adopt the ChatCompletion API of the "gpt-4-turbo-2024-04-09" model to generate QA pairs on the selected subtree:
@@ Appendix D Implementations and Hyperparameters
LABEL:tab:params presents the key training parameters of our primary fine-tuning models: LLaMA-VID, Video LLAVA, and LongVU. Pretraining and implementation details for these models are described below.
LLaMA-VID integrates the EVA-G visual encoder with the QFormer text decoder. Key strategies include freezing both the visual encoder and the text decoder during all stages, except for the uninitialized BERT module, which is optimized for one epoch per stage.
Video-LLaVA processes input images by resizing and cropping them to 224×224, with 8 frames sampled per video. Training is conducted in two stages: the first stage employs a batch size of 256, while the second stage reduces it to 128.
LongVU utilizes SigLIP and DINOv2 for visual encoding, combined with Qwen2-7B for language processing. Training involves cross-entropy loss for text generation, optimized using the AdamW optimizer with a cosine schedule. Image-language pretraining is performed with a batch size of 128 and a learning rate of 1e-5. During video-language fine-tuning, the batch size is reduced to 64, with adaptive compression techniques applied. The model employs sliding windows and spatio-temporal compression methods for efficient processing.
Table 4: Fine-tuning Hyperparameters.  This table details the primary training parameters utilized for fine-tuning LLaMA-VID, Video-LLaVA, and LongVU. Parameters not explicitly listed here adhere to the default configurations of their respective implementations.
Parameter Value
@@
$Pretrained~{}LLM$ $Vicuna\_1.5\_7B$
$max\_length$ $2048$
$batch\_size$ $1$
$gradient\_accumulation$ $64$
$learning\_rate$ $5e^{-6}$
$weight\_decay$ 00
$warmup\_ratio$ $0.03$
$lr\_scheduler$ $cosine$
$image\_token$ $2$
$video\_fps$ $1$
@@
$Pretrained~{}LLM$ $LLaMA\_2\_7B$
$max\_length$ $2048$
$batch\_size$ $32$
$gradient\_accumulation$ $2$
$learning\_rate$ $1e^{-5}$
$weight\_decay$ 00
$warmup\_ratio$ $0.03$
$lr\_scheduler$ $cosine$
$image\_token$ $256$
$video\_frames$ $8$
@@
$Pretrained~{}LLM$ $Qwen\_2\_7B$
$max\_length$ $8192$
$batch\_size$ $32$
$gradient\_accumulation$ $1$
$learning\_rate$ $5e^{-7}$
$weight\_decay$ 00
$warmup\_ratio$ $0.03$
$lr\_scheduler$ $cosine$
$image\_token$ $144$
$video\_fps$ $1$
Appendix E More LongViTU Examples
Table 5: Examples of LongViTU.  We demonstrate the ratio of each concept category and more examples of LongViTU. For a more intuitive perspective, you may refer to Figure 3.
Category Ratio Question Example (Q) Answer Example (A)
Object 12.2% What am I holding in my hand? A cup.
What items are on the table? A pen and a notebook.
Attribute 10.7% What is the color of that clothing? Blue.
What is the material of the cup? Porcelain.
Location 15.5% Where am I in the house right now? In the kitchen.
Where is the key placed in? Inside the drawer.
Action 16.6% What is that man doing? He is cooking.
What am I doing by the counter? Preparing coffee.
Transition 8.1% Where did he go after leaving here? To the garden.
What change happened to that cup? It fell and broke.
Interaction 3.4% Which hand did I use to pick up this wrench? The left hand.
What did I take out of the microwave? A bowl of soup.
Causality 5.4% What happened after I pressed that button? The light turned on.
What happened after I opened the box? A toy popped out.
Motivation 7.5% Why should I leave the room? There is smoke.
Why does she want to open the cabinet? To get a plate.
Planning 5.4% How do I get to the backyard? Through the kitchen door.
How do I repair this house? By fixing the roof first.
Risk 2.7% What dangers does that saw pose? It could cut your hand.
What dangers are there in the kitchen? Slippery floor.
Function 6.4% What is the function of this tool? To tighten bolts.
What is the function of this box? To store items.
Affordance 4.6% What can this stone be used for? To grind spices.
What can this glass bottle be used for? To store water.

### v2

@@ 3 Experiments
Frame-Based Video-LLaVA 45.9 37.8 46.3 49.1 38.1 42.7
@@
Frame-Based Video-LLaVA 45.9 45.6 50.5 48.8 53.2 49.4
@@
Frame-Based Video-LLaVA 45.9 41.6 56.8 55.3 54.6 51.7
@@
LongVU 67.6 55.6 66.0 54.1 46.6 35.7 48.3 51.1 42.8
LongVU∗∗ 70.8 55.9 66.2 54.3 47.0 36.5 48.9 51.4 44.2
@@
LLaVA-Video 57.3 61.9 74.0 60.2 51.6 - - - -
LLaVA-Video∗∗ 62.8 62.2 73.9 60.4 52.1 - - - -
[…]
Method MLVU
Count Ego Needle Order PlotQA Anomaly Reco. Topic Reasoning Avg.
LongVU 27.9 55.7 73.5 53.3 68.5 74.5 86.7 62.9
LongVU∗∗ 29.8 62.5 75.5 54.8 68.8 73.5 86.7 65.2
[…]
LLaVA-Video 38.5 58.2 75.5 65.3 77.0 74.0 85.2 67.7
LLaVA-Video∗∗ 39.4 63.4 76.1 66.8 77.0 74.0 86.7 69.9
[…]
Method LVBench
Key Info. Entity Recog. Event Underst. Reason. Temp. Ground. Summar. Avg.
LongVU 38.1 38.4 34.2 43.5 27.7 28.6 35.1
LongVU∗∗ 38.1 39.8 35.5 44.2 28.8 26.8 35.3
[…]
LLaVA-Video 45.2 42.3 41.4 48.3 33.8 26.8 39.6
LLaVA-Video∗∗ 45.2 42.7 41.2 49.0 35.4 26.8 40.0
[…]
Method MVBench
Action Pred. Episodic Reason. Moving Count Moving Dir. Object Shuffle Unexp. Action
LongVU 66.5 52.5 79.0 90.0 38.5 72.5
LongVU∗∗ 67.0 55.5 80.0 90.5 39.5 77.5
[…]
3.1 Setup
Settings and Models.  We consider a variety of video understanding models, including frame-sampling-based, long-context (therefore, no need to do frame sampling), and multimodal agents. For frame-sampling-based models such as Video-LLaVA [31], we uniformly sample 8 frames per video as input. For long-context LLaMA-VID [30], LongVA-DPO [63], LongVU [43], LLaVA-Video [64] and Gemini-1.5-Pro [41], we sample 1 frame per second. For multi-modal agent, we evaluate VideoAgent [15] and sample 1 frame every 2 seconds. All models (except for VideoAgent) are with an approximate total parameter size of 7B.
Metrics and Benchmarks.  Since LongViTU answers are open-ended, for the evaluations on LongViTU test set, we designed a multi-level scoring system leveraging GPT-4’s near-human comprehension to evaluate answer-ground truth alignment. This approach rewards concise, accurate responses while penalizing hallucinations or irrelevant answers. Detailed criteria and prompts are in subsection C.4. Besides the LongViTU test set, we consider a collection of long-video understanding benchmarks. including ID benchmark EgoSchema [36], as well as OOD benchmarks VideoMME [16], MLVU [66], and LVBench [49]. Additionally, we also loop in canonical video understanding benchmarks (still OOD in terms of video sources), including MVBench [29], WorldQA [65], and OpenEQA [35].
3.2 Quantitative Analysis on LongViTU Test Set
The quantitative results on LongViTU detailed in Table 2 reveal the following critical insights:
Challenges Posed by LongViTU.  Zero-shot evaluation of mainstream models on LongViTU reveals substantial challenges. Even after supervised fine-tuning (SFT), the best-performing model, LongVU, achieves only a GPT-4 score of 55.9, which is far below human performance at 81.0, underscoring a persistent gap. Additionally, proprietary models such as Gemini-1.5-Pro struggle with long-video comprehension, primarily because they must process full-length videos ranging from tens of minutes to hours. Moreover, these models exhibit a pronounced reliance on textual cues rather than spatiotemporal understanding, highlighting a fundamental limitation of current video-language models in handling long-video context and condensed reasoning challenges.
3.3 Quantitative Analysis on Other Benchmarks
To evaluate the impact of SFT on LongViTU, we train several state-of-the-art video understanding models and assess their performance on both ID and OOD scenarios, as detailed in Table 3. Key findings are outlined below.
Broad Effectiveness of SFT.  Fine-tuning on LongViTU leads to substantial performance improvements across nearly all ID, long video OOD, and non-long video OOD benchmarks for major open-source video understanding models. This indicates that the challenges posed by LongViTU (as outlined in Table 2) do not primarily arise from dataset bias, a claim further supported by human-level performance. Instead, LongViTU provides a wealth of diverse knowledge (long video context and condensed reasoning, discussed in section 1), equipping models with enhanced capabilities for video understanding across a broad range of scenarios.
Larger Gains on Longer Videos.  Fine-tuning with LongViTU demonstrates greater improvements for longer videos, as evident from performance trends across different subsets. While results declined on shorter videos, such as the VideoMME Short subset and OpenEQA, where the latter’s average duration (49s) is even shorter than VideoMME Short (83s), significant gains were observed in the Medium (563s) and Long (2386s) subsets. Notably, LLaVA-Video SFT improved by 1% on VideoMME Long, whereas performance slightly decreased on Short (-0.1%) and remained stable on Medium (+0.3%). This pattern is consistent with broader trends in LongVU and LLaVA-Video, where the advantages of SFT become increasingly pronounced as video length increases. These results highlight LongViTU’s effectiveness in enhancing long video context comprehension, reinforcing the benefits of models optimized for extended temporal contexts despite minor trade-offs on shorter videos.
Task-Specific Enhancements.  Both LongVU and LLaVA-Video achieved significant gains in MLVU, primarily driven by improvements in the "Ego" task. These models also demonstrated strong capabilities in "Count" and "Order" tasks, though their performance in other areas remained relatively weaker. On LVBench, where most subcategories do not overlap, the most pronounced improvement was observed in "Temporal Grounding". A similar trend was evident in MVBench, where categories aligned with LongViTU’s predefined categories (notably in condensed reasoning tasks), such as "Episodic Reasoning" and "Unexpected Action", demonstrated significant gains, whereas improvements in other categories were more modest.
3.4 Ablation Study on LongViTU
Due to space constraints, additional details on the ablation study and qualitative evaluation are provided in Appendix B. These analyses offer deeper insights into LongViTU’s adaptability to varying input lengths and its capacity to leverage extended training, highlighting its robustness and scalability.
@@ 4 Conclusion
We’ve introduced LongViTU, a large-scale, automatically generated dataset for long-form video understanding, featuring long video context and condensed reasoning. Fine-tuning canonical VLMs on LongViTU yields substantial performance gains across both ID and diverse OOD long video understanding benchmarks, underscoring its effectiveness.
@@
Appendix A Limitation Statement and Future Works
Limitation Statement.  Our automated pipeline exhibits adaptability across diverse scenarios, self-revision mechanism mitigates QA pairs with pronounced textual biases. However, as illustrated in Figure 4, human evaluation indicates that approximately 9% of QA pairs still retain such biases. This limitation primarily arises from the reliance on LLM-generated multimodal data. Nevertheless, recent studies on vision-language models suggest that a controlled presence of such biases may enhance model robustness.
Future Works.  Future research will focus on enhancing data quality and exploring novel model architectures, e.g., learnable token compression and memory mechanism, etc.
@@ Appendix B Additional Results
Due to space limitations, we defer more details about the ablation study and qualitative evaluation to this section.
@@ B.1 Ablation Study
Varying Durations.  Figure 5 illustrates LongVU’s performance on LongViTU with varying QA durations. As the QA duration (in seconds) increases, performance steadily declines, aligning with the trend observed in VideoMME across all baseline models (Table 3). This downward trend persists across the Short, Medium, and Long subsets in both the zero-shot and SFT settings. This result is intuitive, as longer videos naturally introduce greater complexity and challenge compared to shorter ones.
Figure 5: Duration Ablation.  Regardless of whether in the zero-shot or fine-tuned setting on LongViTU, LongVU demonstrates a performance decline on longer subsets [300, 600) compared to shorter ones [60, 300) and [0, 60).
Training Epochs.  Figure 6 illustrates that LongVU’s performance on Egoschema improves following SFT as the number of training epochs inherited from LongViTU increases. However, the performance gains gradually diminish as the epochs approach the full training scale of LongViTU, eventually stabilizing at its final performance.
@@
B.2 Qualitative Evaluation
To facilitate a more comprehensive qualitative analysis, we provide visualizations of various question-answering scenarios in Figure 7. Due to space constraints, these figures have been deferred to this section.
Spatial Details.  As depicted in Figure 7(a), the dense arrangement of multiple foreground objects in the scene led to incorrect zero-shot predictions by both Video-LLaVA and LongVU. Fine-tuning with LongViTU enabled the model to better capture fine-grained spatial features, resulting in correct answers.
Key Moments.  As illustrated in Figure 7(b), the fine-tuned models successfully identified the key moment, a brief appearance of "a plant on the windowsill", and provided a precise and concise response, achieving a perfect GPT-4 score of 100. In contrast, the Video-LLaVA zero-shot model failed to capture this brief key moment.
Temporal Localization.  As shown in Figure 7(c), both Video-LLaVA and LongVU accurately identified "two" plug-in sockets in the kitchen at the end of a long video sequence following fine-tuning. However, in the zero-shot setting, both models struggled with this task. This observation underscores the challenge of extracting spatial details from extended video sequences and highlights the contribution of LongViTU in enhancing generalization for long-form temporal localization.
(a) Spatial Details.  The dense arrangement of multiple foreground objects in the scene led to incorrect zero-shot predictions by both Video-LLaVA and LongVU. Fine-tuning with LongViTU enabled the model to better capture fine-grained spatial features, resulting in correct answers.
(b) Key Moments.  The fine-tuned models successfully identified the key moment, a brief appearance of "a plant on the windowsill", and provided a precise and concise response, achieving a perfect GPT-4 score of 100. In contrast, the Video-LLaVA zero-shot model failed to capture this brief key moment.
(c) Temporal Localization.  Both Video-LLaVA and LongVU accurately identified "two" plug-in sockets in the kitchen at the end of a long video sequence following fine-tuning. However, in the zero-shot setting, both models struggled with this task. This observation underscores the challenge of extracting spatial details from extended video sequences and highlights the contribution of LongViTU in enhancing generalization for long-form temporal localization.
@@ C.2 Long-Form QA Generation
QA Generation.  We adopt the ChatCompletion API of the "gpt-4-turbo-2024-04-09" model to generate QA pairs on the selected sub-tree:
@@ C.4 Evaluation Metrics
C.5 Category Definitions
To enhance Large Language Models (LLMs) in constructing episodic memory from video content, we employ a structured taxonomy for question generation. As described in subsubsection 2.1.2, the LLM assigns each generated question to an appropriate category within this taxonomy, ensuring more condensed and targeted reasoning. Below, we provide the specific definitions and examples for each category.
  • $\bullet$
    Object:  Questions identifying or referencing physical entities present in the video.
    Question: What was the red object on the table?
    Answer: A red coffee mug.
  • $\bullet$
    Attribute:  Questions focusing on descriptive properties (e.g., color, shape, texture, size) of objects or entities.
    Question: What was unusual about the person’s jacket?
    Answer: It was bright yellow with reflective stripes.
  • $\bullet$
    Location:  Questions about the spatial placement of objects, characters, or events within the scene.
    Question: Where was the book before it was picked up?
    Answer: On the wooden shelf next to the lamp.
  • $\bullet$
    Action:  Questions regarding physical movements or activities performed by characters or objects.
    Question: What did the man do after opening the door?
    Answer: He walked inside and placed his bag on the table.
  • $\bullet$
    Transition:  Questions involving contrasts or similarities between objects, actions, or scenes.
    Question: How did the first explosion differ from the second?
    Answer: The first was small and controlled, while the second was much louder and caused debris to scatter.
  • $\bullet$
    Interaction:  Questions that explore connections between entities, including people, objects, or events.
    Question: How was the old man related to the boy?
    Answer: He was the boy’s grandfather.
  • $\bullet$
    Causality:  Questions about cause-and-effect relationships within events.
    Question: Why did the lights go out suddenly?
    Answer: Because a power surge occurred.
  • $\bullet$
    Motivation:  Questions exploring the intent or reasoning behind a character’s actions.
    Question: Why did the thief hesitate before taking the wallet?
    Answer: He noticed a security camera watching him.
  • $\bullet$
    Planning:  Questions related to future intentions or strategies within the video.
    Question: What was the team’s plan to escape?
    Answer: They decided to climb through the air ducts.
  • $\bullet$
    Risk:  Questions that examine potential danger, hazards, or consequences of an action.
    Question: What was the risk of crossing the bridge?
    Answer: It was old and unstable, with missing planks.
  • $\bullet$
    Function:  Questions addressing the purpose or intended use of an object or action.
    Question: Why did the character press the button?
    Answer: To call the elevator.
  • $\bullet$
    Affordance:  Questions about how an object can be used or interacted with, based on its design or properties.
    Question: What did the woman use the chair for?
    Answer: She stood on it to reach the top shelf.
@@ Appendix D Implementations and Hyperparameters
The fine-tuning process for open-source models is governed by a set of key hyperparameters that influence their performance and generalization capabilities. Table 4 provides a structured overview of these critical training parameters for LLaMA-VID [30], Video-LLaVA [31], LongVU [43], and LLaVA-Video [64]. Each model has been fine-tuned with distinct configurations to optimize its ability to process and understand video-based inputs effectively. Details regarding the pretraining strategies and implementation specifics of these models are anonymously available at https://longvitu.github.io.
Table 4: Fine-tuning Hyperparameters.  This table presents the essential hyperparameters employed during fine-tuning for the latest mainstream open-source models, all of which have 7B parameters for a fair comparison. Parameters that are not explicitly specified adhere to the default configurations of the respective model implementations.
Hyperparameter Value
@@
Pretrained LLM Vicuna-1.5-7B
Maximum sequence length 2048
Batch size 1
Gradient accumulation steps 64
Learning rate $5e^{-6}$
Weight decay 0
Warmup ratio 0.03
Learning rate scheduler Cosine
Image tokens per sample 2
Video frame rate (FPS) 1
@@
Pretrained LLM LLaMA-2-7B
Maximum sequence length 2048
Batch size 32
Gradient accumulation steps 2
Learning rate $1e^{-5}$
Weight decay 0
Warmup ratio 0.03
Learning rate scheduler Cosine
Image tokens per sample 256
Video frames per sample 8
@@
Pretrained LLM Qwen-2-7B
Maximum sequence length 8192
Batch size 32
Gradient accumulation steps 1
Learning rate $5e^{-7}$
Weight decay 0
Warmup ratio 0.03
Learning rate scheduler Cosine
Image tokens per sample 144
Video frame rate (FPS) 1
LLaVA-Video
Pretrained LLM Qwen-2-7B
Maximum sequence length 32768
Batch size 1
Gradient accumulation steps 1
Learning rate $1e^{-7}$
Vision tower learning rate $2e^{-8}$
Weight decay 0
Warmup ratio 0.03
Learning rate scheduler Cosine
Image tokens per sample 196
Video frames per sample 64
Appendix E Related Work
Large Language Models.  Large language models (LLMs), including InstructGPT [37], GPT-4 [1], LLaMA [47], and LLaMA-2 [48], have demonstrated remarkable capabilities in text processing and large-scale dataset generation. These models convert multimodal inputs into structured textual representations, which can then be used to prompt GPT-4 for multimodal data synthesis.
Instruction Tuning.  LLaVA [33] pioneered the use of foundational vision models to generate image captions and detect bounding boxes, which were subsequently processed by ChatGPT or GPT-4 to create image-based QA datasets. Building upon this, approaches such as Bongard-OpenWorld [52], Video-LLaVA [31], and VideoChat [28] extended these techniques to video-based QA. By sampling multiple frames from videos and applying LLaVA’s methodology, these methods generate structured video QA datasets using frame-wise descriptions, object categories, and attributes. However, the redundancy in frame-based descriptions, coupled with the input length constraints of LLMs, limits the number of frames that can be processed, thereby reducing dataset comprehensiveness.
Long-Context Language Models.  Despite advances in long-context LLMs [45], including GPT-4-turbo, ChatGLM [18], Baichuan2 [56], and InternLM2 [6], these models exhibit significant performance degradation when processing extended and intricate texts. They struggle to manage the redundancy and structural disorganization inherent in detailed video descriptions, limiting their effectiveness in generating video QA. Unlike static images, video understanding requires modeling temporal dynamics, making event correlation essential. However, existing frame-based approaches fail to capture this aspect adequately, often resulting in shallow QA generation confined to frame-level analysis.
Long-Form Video Understanding.  Instruction tuning within the LLaVA paradigm [33] has proven effective for multimodal tasks such as captioning [8] and visual question answering (VQA) [4, 3, 46, 27, 12, 57, 2, 20, 38, 51]. While these methods perform well on images and short videos, they face significant challenges when applied to long-form videos [44, 31, 34, 61, 49, 63]. A primary bottleneck arises from the large number of visual tokens produced by encoders—ranging from 576 to 2880 tokens per image in models like LLaVA-NeXT [32]—which quickly exceed the context window limits of LLMs as the number of frames increases. To address this, recent methods employ resampling techniques to reduce visual tokens before feeding them into LLMs [27, 30, 5, 10]. However, such reductions often degrade visual representations, leading to performance deterioration. Alternative strategies, including advanced token pruning and feature merging, offer promising solutions [9, 42, 23, 67]. Notably, recent models such as VideoLLM Online [7], LongVILA [55], LLaVA-OV [26], LongVA [63], and LongVU [43] represent significant advancements in addressing these limitations and enhancing long-form video understanding.
Appendix F More LongViTU Examples
The following presents more examples of the detailed categorization of LongViTU.
(a) Spatiotemporal Understanding.  This subfigure presents examples for four subcategories within Spatialtemporal Understanding, including Object, Attribute, Location, Action.
(b) Episodic Reasoning.  This subfigure presents examples for four subcategories within Episodic Reasoning, including Transition, Interaction, Causality, Motivation.
(c) Commonsense Inference.This subfigure presents examples for four subcategories within Commonsense Inference, including Planning, Risk, Function, Affordance.
Figure 8: LongViTU Examples.  For better illustration, only concise key information is presented in the predictions.